OpenStudy (anonymous):

Suppose that X and Y have the follolwing joint pdf: f(x,y)=c(x2+y) for 0≤y≤1−x^2 and f(x,y)=0 for otherwise. Determine the value of constant,c?

OpenStudy (anonymous):

the question got mistake.. f(x,y) = c(x^2+y)

OpenStudy (anonymous):

yeah this is going to amount so setting some damn double integral equal to 1 and solving for c i have to go, and probably would make a mistake anyway, but try mimicing probem 1 on page 7 here

OpenStudy (anonymous):

I know we need to use double integral equal to 1...i confuse what is the domain for the integral?

OpenStudy (anonymous):

\[\int\limits_{a}^{b} \int\limits_{c}^{d} f(x,y) dx dy =1\]

OpenStudy (anonymous):

what is a, b , c, and d for equation above?

OpenStudy (turingtest):

f(x,y)=c(x^2+y) for 0≤y≤1−x^2 ^^^^ these are the bounds on y and f(x,y)=0 otherwise ^^^if this is true, what are the bounds on x?

OpenStudy (zarkon):

you already asked this question http://openstudy.com/users/zarkon#/updates/4f9ac288e4b000ae9ed03916